OpenStudy (anonymous):

Easy one? Need help! "If you roll 3 dices, what is the probability that their sum will be 9?"

OpenStudy (anonymous):

How do I go about figuring this one out?

OpenStudy (anonymous):

I found this online.. but im not for sure that its all of the probabilities. 9 = 6 + 2 + 1 = 4 + 3 + 2 = 3 + 3 + 3 = 2 + 2 + 5 = 1 + 3 + 5 = 1 + 4 + 4

OpenStudy (anonymous):

Okay, so those are the total numbers of fortunate outcomes. How do I know the number of possible outcomes? 6^3? Why?

OpenStudy (anonymous):

Any know how I can calculate the total number of possible outcomes?

OpenStudy (anonymous):

According to http://statistics.about.com/od/ProbHelpandTutorials/a/Probabilities-For-Rolling-Three-Dice.htm there are 25 possible outcomes, and so the probability is 11.6%

OpenStudy (anonymous):

But they only list 6 different possibilites of a sum of 9. Anyone can explain?

Directrix (directrix):

Die: a small cube marked on each face with from one to six spots. There are 25 tosses that result in a sum of 9. Think of the three die as being different colors so that you can differentiate among them. 126, 135, 144, 153, 162; 216, 225, 234, 243, 252, 261; 315, 324, 333, 342, 351; 414, 423, 432, 441; 513, 522, 531; and 612, 621. The three die can be cast in 6*6*6 = 216 outcomes. There are 6 outcomes for the first die, 6 for the second, and 6 for the third. 25/216 is the probability of tossing a sum of 9 in one toss of three die. @Veritas45